OpenStudy (anonymous):

Integrate : ∫ (xcosx)/(xsinx + cosx)^2 dx

OpenStudy (anonymous):

its (xsin(x)+cos(x))^2?

OpenStudy (anonymous):

if its like that then try substitude u=xsin(x)+cos(x)

OpenStudy (anonymous):

\[\int\limits \frac{xcosx}{(xsinx+cosx)^2} dx\] I'm sorry .. That's the correct question.

OpenStudy (anonymous):

ok then let u=xsin(x)+cos(x) and solve

OpenStudy (anonymous):

Done. Then du = xcosx-sinx

OpenStudy (anonymous):

um no u have some mistake there

OpenStudy (anonymous):

Oh!

OpenStudy (anonymous):

ooooooooo!! Thanks ! I got it!

OpenStudy (anonymous):

easy right ? :) , ur welcomed

OpenStudy (anonymous):

:)